The area of a rectangle is A=LW, where the L is the length of the rectangle and W is the width. Find the area of the rectangle if it's length is 9 feet and its width is 14 feet less than twice it's length.
Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39617)

Length, 9
Width, W=2*9-14=4

Multiply 9 and 4.

Answer by MathTherapy(10552)

The area of a rectangle is A=LW, where the L is the length of the rectangle and W is the width. Find the area of the rectangle if it's length is 9 feet and its width is 14 feet less than twice it's length.
With length being 9 feet, width = 2(9) - 14 = 4 feet
Correct answer:
FYI: It's NOT 9(18).